# http://integratedcognition.ai

### v1

@@ Abstract
One goal of AI (and AGI) is to identify and understand specific mechanisms and representations sufficient for general intelligence. Often, this work manifests in research focused on architectures and many cognitive architectures have been explored in AI/AGI. However, different research groups and even different research traditions have somewhat independently identified similar/common patterns of processes and representations or cognitive design patterns that are manifest in existing architectures. Today, AI systems exploiting large language models (LLMs) offer a relatively new combination of mechanism and representation available for exploring the possibilities of general intelligence. In this paper, we summarize a few recurring cognitive design patterns that have appeared in various pre-transformer AI architectures. We then explore how these patterns are evident in systems using LLMs, especially for reasoning and interactive (“agentic”) use cases. By examining and applying these recurring patterns, we can also predict gaps or deficiencies in today’s LLM systems and identify likely subjects of future research towards general intelligence using LLMs and other generative foundation models.
@@ 1 Introduction
Cognitive architectures encapsulate and represent theories and commitments toward a general, systems architecture for intelligence. More than one hundred different cognitive architectures have been proposed and developed, drawing from many distinct (and often quite disparate) intellectual traditions [15]. Remarkably, as architectures have evolved in response to research outcomes, even though they draw from very different sources, there has been notable convergence and even consensus around both a high-level functional architecture of cognition [18] as well as many lower-level algorithmic [13] and representational [5, 6] commitments. A tentative but hopeful inference from such convergence is that, as a field, we are beginning to understand what components are necessary (or at least important) for realizing artificial general intelligence (AGI).
In contrast to the comparatively long history of cognitive architecture, large language and multi-modal models (LLMs [^1] Throughout, we use LLM as the common term used to describe these models, although most recent models have input and output modalities beyond text/language.) have emerged in the past five years as a major new technology, capturing significant research interest, investment, and even popular attention. The attention is motivated by the breadth and depth of these models and their applicability to many different use cases and applications. While it is sometimes claimed that LLMs alone might offer the potential for AGI [3], a new and rapid growing research area is directed toward exploiting LLMs as one (important) component within a larger collection of components and tools that comprise a general AI system [14, 63, 24].
In this paper, we focus on systems architectures comprised primarily of LLM components (as opposed to more hybrid integrations of LLMs with planners [14], constraint solvers [22], traditional cognitive architectures [17], etc.). Such Agentic LLM Systems often are developed for specific tasks or domains, such as software development [8] or web-content publishing [36]. However, researchers and developers are also attempting to identify and to develop agentic frameworks sufficiently general to apply to any task [62, 37, 55]. Further, Agentic LLM Systems are often enhanced with specialized memories that are also common in cognitive architecture (e.g., episodic memory). In sum, the search for general, Agentic LLM Systems shares many goals and non-trivial overlap with cognitive architecture research.
@@ 2 Cognitive Design Patterns
In software engineering, a design pattern is a description of or template for a solution to a particular design problem. Design patterns are typically specified abstractly (not code) and are flexible enough to be adapted/customized for specific situations. Cognitive design patterns represent a similar concept. A cognitive design pattern denotes some function/process or representation/memory that routinely appears in agents and/or cognitive architectures. That is, like software design patterns, cognitive design patterns summarize, at an abstract level, common (or at least recurring) solutions to functional requirements for intelligent systems. Table 1 lists some examples of cognitive design patterns, drawn from past comparative analyses from various researchers [13, 18, 15, 38]. For illustration, we include examples from the ACT-R [1] and Soar [16] cognitive architectures and from the belief-desire-intention (BDI) family of agent architectures [53]. We introduce cognitive design patterns to support the organization and unification of such analyses.
@@
Importantly, cognitive design patterns, like software design patterns, are abstract. They not only abstract implementation details, but even specific algorithms. As suggested by examples in the table, distinct algorithms and representational commitments are employed to implement a cognitive design pattern in an individual architecture. An example is illustrated in Figure 1. Many architectures use a 3-stage process to control asserting and retracting individual memory items (rather than the more common store/erase, 2-stage process). Instead of direct assertion, an agent generates candidates for some memory slot. It then uses a selection or commitment process [53] to choose among the candidates before making a final choice or intention.
@@
Familiar algorithms may be specialized (in scope and/or implementation) to realize a cognitive function within an architecture. For example, Soar uses a justification-based truth maintenance system (JTMS) [9] as the reconsideration process for one functional type of memory element. The JTMS implementation is specialized for its integration within Soar. Cognitive design patterns are useful as a unit of comparison and analysis because abstracting to just identifying a functional role (reconsideration) distances the analysis from implementation details (a customized JTMS for Soar), making more evident and clear commonalities and contrasts in architectural commitments. If we just assert that Soar uses a JTMS and a BDI architecture uses some decision-theoretic calculations over different types of memory instances, it is easy to miss that the functional role of both of these processes, reconsideration (and more broadly, belief revision and context/memory management), is comparable in both architectures.
@@ 3 Cognitive Design Patterns & Agentic LLM Systems
    Are there unique properties of Agentic LLMs that suggest new cognitive design patterns for general intelligence? We investigate whether the computational and behavioral characteristics of Agentic LLMs suggest new cognitive design patterns.
The following subsections consider each of these questions, drawing examples from research in both Agentic LLMs and cognitive and agent architectures. We identify specific examples to illustrate and (hopefully) provide insight. An exhaustive survey for each of these questions is too broad for a conference paper and, given the accelerating exploration of these topics, somewhat impractical.
@@ 3.1.1 Observe-Decide-Act in ReAct
Generally, LLM performance improves when the user prompt directs the model to be explicit about its reasoning, as in the now ubiquitous Chain-of-Thought (CoT) prompting method [52]. ReAct [62], one of the earliest and most influential Agentic LLM methods, builds on CoT by distinguishing between internal and external problem-solving steps. As illustrated in Figure 2 (left), ReAct initially prompts the LLM for natural-language (NL) statements (thoughts) about the problem before taking any action. Directing the LLM to separate its reasoning and action steps results in significant improvements in LLM responses across multiple domains [62]. ReAct is now a foundational approach for Agentic LLM systems and has been incorporated into agentic software development tools such as LangGraph [21].
@@
ReAct has a significant overlap with this pattern but lacks the step that explicitly makes commitments. Thus, this mapping of ReAct to the familiar design pattern leads to an immediate, empirical question: Would introducing a notion of commitment in ReAct result in better overall reasoning outcomes than ReAct alone? We contend that one of the advantages of mapping Agentic LLM solutions to cognitive design patterns is identifying specific questions such as this one.
@@ 3.1.2 Episodic Memory in Generative Agents
One early, influential example of using long-term memory in Agentic LLM Systems is Generative Agents [31]. In this research system focused on human behavior simulation, an LLM agent represents a person living and working in a small village. Different agents interact with one another through natural language and undertake various individual and collective tasks, such as attending work or school, running for office, planning a party, etc. Generative Agents implements a long-term memory that concatenates a history of specific recent experiences (observations) and abstract, integrative summaries of past experience (reflections). Reflections capture gists of experience (“Klaus worked in the lab all day”) and also insights (“Klaus’ dedication to his lab work suggests he is passionate about research.”).
These memories of past experiences then inform the subsequent behavior of that agent. Agents retrieve a combination of observations and reflections into a Memory Stream, which provides the context for agent planning and action. Retrieval uses recency, agent-judged importance (at the time of memory encoding), and an agent’s subjective sense of relevance to retrieve items from long-term memory into the memory stream.
@@
The developers of Generative Agents did not describe its long-term memory as an episodic memory. However, this paper is often cited [^3] Over 1,500 total citations according to Semantic Scholar. as an example of episodic memory using LLMs. Episodic memory is autobiographical, a memory of what happened to me. Specific characteristics of human episodic memory are well-understood [44] and have been researched and implemented within cognitive architectures [28].
@@
To further highlight one example from the table, most conceptions of episodic memory assume the encoding specificity principle [45], which (roughly) says that episodic memories encode the context of events. Retrieval of some past episode is more likely when retrieval cues match the encoded context: entering a home that one lived in many years ago might trigger memories of being in that home. The specific context (e.g., the floor plan) was encoded in memory. Being present in the home again creates perceptual cues that match encoded cues from past memories, resulting in retrieval.
The memory in Generative Agents encodes objects, providing some specificity, but the encoding of observations and reflections is not fully contextual. Retrieval instead employs relevance, which relies on semantic similarity rather than cue-specificity. The table comments on other similarities and differences as well. As for ReAct, this analysis points to specific research questions, such as exploring how agent behavior might change if the agent could deliberately attempt to construct explicit retrieval cues or whether the quality of reflection would change with a more context-sensitive trigger for reflection.
@@ 3.2 Patterns Apt for Exploration and Exploitation
In this section, we briefly discuss two cognitive design patterns that have not yet been taken up in the Agentic LLM mainstream: reconsideration and knowledge compilation. As detailed further, both patterns appear to be responsive to current shortcomings in Agentic LLMs and thus represent research opportunities.
@@ 3.2.1 Commitment and Reconsideration in Agentic LLMs
Chat-focused LLMs tend to resist major redirection in the trajectory of their token generation. Further, the ability of LLMs alone to produce non-monotonic reasoning steps has demonstrated somewhat mixed results [23]. In Agentic LLMs, various approaches have been developed, often comparable to short-term memories, to limit or to manage what context is presented with each LLM query [30], comparable to paging in operating systems. This approach does support (if somewhat indirectly) redirection by forgetting/deleting past context.
As above, reconsideration is the cognitive design pattern that encapsulates the process of evaluating whether a prior commitment, such as a goal or a plan, should continue given the agent’s current situation. When reconsideration results in a change in commitment, the direction of agent reasoning changes non-monotonically. In addition to general reflection strategies (discussed below), the use of reconsideration for intentions or commitments in Agentic LLMs could allow an the agent to periodically assess the continuing viability and desirability of its goals. This research direction would enhance recent efforts in using Agent LLMs for planning and plan decomposition, which share some similarities to commitment processes [33, 60]. As LLM Agents make deliberate, explicit commitments, they will then also need to decide if/when those commitments should be abandoned, just as traditional agents do.
@@ 3.2.2 Knowledge Compilation
Multi-step reasoning (more generally, problem search [19]) can be computationally expensive, even NP-complete [4]. Knowledge compilation caches the results of (possibly many) reasoning steps into a more compact representation, essentially amortizing the cost of expensive reasoning by enabling the immediate use of the cached steps in future, similar situations. Many algorithms have been developed for knowledge compilation, including macro-operator learning in STRIPS [10] and various forms of explanation-based learning [26]. Knowledge compilation is a core learning algorithm in both ACT-R (production compilation [42]) and Soar (chunking [20]).

### v2

@@ Abstract
One goal of AI (and AGI) is to identify and understand specific mechanisms and representations sufficient for general intelligence. Often, this work manifests in research focused on architectures and many cognitive architectures have been explored in AI/AGI. However, different research groups and even different research traditions have somewhat independently identified similar/common patterns of processes and representations or cognitive design patterns that are manifest in existing architectures. Today, AI systems exploiting large language models (LLMs) offer a relatively new combination of mechanisms and representations available for exploring the possibilities of general intelligence. This paper outlines a few recurring cognitive design patterns that have appeared in various pre-transformer AI architectures. We then explore how these patterns are evident in systems using LLMs, especially for reasoning and interactive (“agentic”) use cases. By examining and applying these recurring patterns, enables predictions of gaps or deficiencies in today’s Agentic LLM Systems and identification of subjects of future research towards general intelligence using generative foundation models.
@@ 1 Introduction
Cognitive architectures encapsulate and represent theories and commitments toward a general systems architecture for intelligence. More than one hundred different cognitive architectures have been proposed and developed, drawing from many distinct (and often quite disparate) intellectual traditions [17]. Remarkably, as architectures have evolved in response to research outcomes, even though they draw from very different sources, there has been notable convergence and even consensus around both a high-level functional architecture of cognition [20] as well as many lower-level algorithmic [15] and representational [6, 7] commitments. A tentative but hopeful inference from such convergence is that, as a field, we are beginning to understand what components are necessary (or at least important) for realizing artificial general intelligence (AGI).
In contrast to the comparatively long history of cognitive architecture, large language and multi-modal models (LLMs [^1] Throughout, we use LLM as the common term used to describe these models, although most recent models have input and output modalities beyond text/language.) have emerged in the past five years as a major new technology. The attention is motivated by the breadth and depth of these models and their applicability to many different use cases and applications. While it is sometimes claimed that LLMs alone might offer the potential for AGI [4], a new and rapid growing research area is directed toward exploiting LLMs as one (important) component within a larger collection of components and tools that comprise a general AI system [16, 67, 26].
Here, we examine systems architectures comprised primarily of LLM components (in contrast to hybrid integrations of LLMs with planners [16], constraint solvers [24], traditional cognitive architectures [19], etc.). Such Agentic LLM Systems often are developed for specific tasks or domains, such as software development [9] or web-content publishing [38]. However, researchers and developers are also attempting to identify and to develop agentic frameworks sufficiently general to apply to any task [66, 39, 59]. Further, Agentic LLM Systems are increasingly enhanced with specialized memories also common in cognitive architecture (e.g., episodic memory). In sum, the search for general Agentic LLM Systems shares many goals and non-trivial overlap with cognitive architecture research.
@@ 2 Cognitive Design Patterns
In software engineering, a design pattern is a description of or template for a solution to a particular design problem. Design patterns are typically specified abstractly (not code) and are flexible enough to be adapted/customized for specific situations. Cognitive design patterns represent a similar concept. A cognitive design pattern denotes some function/process or representation/memory that routinely appears in agents and/or cognitive architectures. That is, like software design patterns, cognitive design patterns summarize, at an abstract level, common (or at least recurring) solutions to functional requirements for intelligent systems. Table 1 lists some examples of cognitive design patterns, drawn from past comparative analyses from various researchers [15, 20, 17, 40]. For illustration, we include examples from the ACT-R [1] and Soar [18] cognitive architectures and from the belief-desire-intention (BDI) family of agent architectures [56]. We introduce cognitive design patterns to support the organization and unification of such analyses. [^2] The coinage here is not new or unique. The the recognition of common patterns occurring across cognitive and agent architectures goes back many years [13] and has been of long-standing interest to the authors [57, 15, 46, 7, 20].
@@
Cognitive design patterns, like software design patterns, are abstract, eliding not only implementation details, but specification of algorithms. Distinct algorithms and representational commitments are employed to instantiate a cognitive design pattern in different architectures. An is illustrated in Figure 1, many architectures use a 3-stage process to control asserting and retracting individual memory items (rather than the more common store/erase, 2-stage process). Instead of direct assertion, an agent generates candidates for some memory slot. It then uses a selection or commitment process [56] to choose among the candidates before making a final choice or intention.
@@
Familiar algorithms may be specialized (in scope and/or implementation) to realize a cognitive function within an architecture. For example, Soar uses a justification-based truth maintenance system (JTMS) [10] as the reconsideration process for one functional type of memory element. The JTMS implementation is specialized for its integration within Soar. Cognitive design patterns are useful as a unit of comparison and analysis because abstracting to just the functional role (reconsideration) distances the analysis from implementation details (a customized JTMS for Soar), making more evident and clear commonalities and contrasts in architectural commitments. If we just assert that Soar uses a JTMS and a BDI architecture uses some decision-theoretic calculations over different types of memory instances, it is easy to miss that the functional role of both of these processes, reconsideration (and more broadly, belief revision and context/memory management) is comparable in both architectures.
@@ 3 Cognitive Design Patterns & Agentic LLM Systems
    Do the unique properties of Agentic LLMs suggest new cognitive design patterns for general intelligence? We examine how computational and behavioral characteristics of Agentic LLMs may lead to new design patterns.
We now consider each of these questions, drawing examples from research in both Agentic LLMs and cognitive and agent architectures. We identify specific examples to illustrate and provide insight. An exhaustive survey for each of these questions is too broad for a conference paper and, given the accelerating pace of exploration of these topics, somewhat impractical.
@@ 3.1.1 Observe-Decide-Act in ReAct
Generally, LLM performance improves when the user prompt directs the model to be explicit about its reasoning, as in the now ubiquitous Chain-of-Thought (CoT) prompting method [55]. ReAct [66], one of the earliest and most influential Agentic LLM methods, builds on CoT by distinguishing between internal and external problem-solving steps. As illustrated in Figure 2 (left), ReAct initially prompts the LLM for natural-language (NL) statements (thoughts) about the problem before taking any action. While subsequent research has interrogated the relative contributions of examples vs. decomposition in ReAct [3], it did demonstrate significant improvements in LLM responses across multiple domains [66]. ReAct is now a foundational approach for Agentic LLM systems and has been incorporated into agentic software development tools such as LangGraph [23].
@@
ReAct has a significant overlap with this pattern but lacks the step that explicitly makes commitments. Thus, this mapping of ReAct to the familiar design pattern leads to an immediate, empirical question: Would introducing commitment in ReAct result in better overall reasoning outcomes than ReAct alone? We contend that one of the advantages of mapping Agentic LLM solutions to cognitive design patterns is identifying specific questions such as this one.
@@ 3.1.2 Episodic Memory in Generative Agents
An early, influential example of LTM in Agentic LLM Systems is Generative Agents [33]. Generative Agents simulates human behavior via a collection of LLM agents, each representing a person living in a small village. Agents interact with one another through natural language and undertake various individual and collective tasks, such as attending work or school. Generative Agents implements an LTM that concatenates a history of specific recent experiences (observations) and abstract, integrative summaries of past experience (reflections). Reflections capture gists of experience (“Klaus worked in the lab all day”) and also insights (“Klaus’ dedication to his lab work suggests he is passionate about research.”).
These memories then inform subsequent agent behavior. Agents retrieve a combination of observations and reflections into a Memory Stream, which provides context for agent planning and action. Retrieval uses recency, agent-judged importance, and an agent’s subjective sense of relevance to retrieve items from long-term memory into the memory stream.
@@
The developers of Generative Agents do not describe its long-term memory as an episodic memory but it is often cited as an example of episodic memory using LLMs. Episodic memory is autobiographical, memory of what happened to me. Specific characteristics of human episodic memory are well-understood [47] and have been implemented within cognitive architectures [30].
@@
To highlight one example, most conceptions of episodic memory assume the encoding specificity principle [48], which (roughly) says that episodic memories encode event context. Retrieval of some past episode is more likely when retrieval cues match the encoded context: entering a home that one lived in many years ago might trigger memories of being in that home. The context (e.g., the floor plan) was encoded in memory. Being present in the home again creates perceptual cues that match encoded cues from past memories, resulting in retrieval.
Memory in Generative Agents encodes objects, providing some specificity, but the encoding of observations and reflections is not fully contextual. Retrieval employs relevance, which relies on semantic similarity rather than cue-specificity. The table identifies other similarities and differences. As for ReAct, this analysis points to specific research questions, such as exploring how agent behavior might change if the agent could deliberately construct explicit retrieval cues or whether the quality of reflection would change with a more context-sensitive trigger.
@@ 3.2 Patterns Apt for Exploration and Exploitation
In this section, we briefly discuss two cognitive design patterns that have not yet been taken up in the Agentic LLM mainstream: reconsideration and knowledge compilation. Both patterns appear to be relevant to addressing current shortcomings in Agentic LLMs and thus represent research opportunities.
@@ 3.2.1 Commitment and Reconsideration in Agentic LLMs
Reconsideration is a cognitive design pattern that describes the process of evaluating whether a prior commitment, such as a goal or a plan, should continue in an agent’s current situation. Reconsideration that results in a change in commitment is a non-monotonic step. Chat-focused LLMs tend to resist major redirection in the trajectory of their token generation and LLMs alone are not consisent or reliable in producing non-monotonic reasoning steps [25].
In addition to reflection strategies (discussed below), reconsideration for intentions or commitments in Agentic LLMs could allow an the agent to periodically assess the continuing viability and desirability of its goals. A few approaches have been developed, often comparable to short-term memories, to limit or to manage what context is presented with each LLM query [32], comparable to paging in operating systems. This approach does support (if somewhat indirectly) redirection by forgetting/deleting past context. Reconsideration via LLMs would enhance recent efforts in using Agent LLMs for planning and plan decomposition, which share some similarities to commitment processes [35, 64]. As LLM Agents make deliberate, explicit commitments, they will then also need to decide if/when those commitments should be abandoned, just as traditional agents do.
@@ 3.2.2 Knowledge Compilation
Multi-step reasoning (or problem search [21]) can be computationally expensive, even NP-complete [5]. Knowledge compilation caches the results of reasoning steps into a more compact representation, essentially amortizing the cost of expensive reasoning by enabling immediate use of the cached reasoning steps in future, similar situations. Examples of knowledge compilation abound, including macro-operator learning in STRIPS [11], various forms of explanation-based learning [28] and implementations in both ACT-R (production compilation [44]) and Soar (chunking [22]).
@@ 4 Conclusions
{credits}
4.0.1 Acknowledgements
This work was partially supported by the Office of Naval Research, contract N00014-21-1-2369, and the Defense Advanced Research Projects Agency, contract HQ072716D0006. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of Defense, Office of Naval Research, or DARPA. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation hereon. Cognitive design patterns have been a decades-long interest of the authors and developed over many interactions and conversations with our colleagues including Ron Chong, Jacob Crossman, Kevin Gluck, Scott Huffman, Randy Jones, Rick Lewis, Christian Lebiere, Doug Pearson, Frank Ritter, Glenn Taylor, Mike van Lent, and Richard Young. We would like to thank the anonymous reviewers for constructive suggestions and feedback.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.